OpenStudy (anonymous):

Use addition or substitution to find the value of x for this set of equations. y = x - 2 7y - 4x = 10

OpenStudy (radar):

Perfect for substitution as first equation has an isolated y in terms of x. so substitute value of y in terms of x in the second equation. You should get this: 7(x-2) -4x =10 clear the parentheses and get: 7x -14 -4x = 10 Now you are home free, just algebra now solve for x.

OpenStudy (radar):

After getting a value for x, substitute in the first equation the value for x.

OpenStudy (anonymous):

whats does x =

OpenStudy (radar):

7x-14-4x=10 Combine the x terms (just as you have in the past) What do you get when 7x -4x is combined?

OpenStudy (anonymous):

is it -3x

OpenStudy (radar):

It is just plain old 3x (positive 7 is larger than -4) Now you have : 3x-14 = 10 add 14 to both left and right of the equal sign. Getting 3x-14 +14 = 10 +14 This now grade school arithmetic 3x=24 See now the x term is on the left and the value is on the right. To solve for x, divide both sides by 3. What is x = to???

OpenStudy (anonymous):

u tell me

OpenStudy (radar):